OpenStudy (anonymous):

What is the simplified form of each expression? (3x^7/2)^6(x^2)^6

OpenStudy (jdoe0001):

\(\large \bf (3x^{\frac{7}{2}})^{\color{red}{ 6}}(x^2)^{\color{red}{ 2}}\implies 3^{\color{red}{ 6}}x^{\frac{7}{2}\cdot {\color{red}{ 6}}}\cdot x^{2\cdot {\color{red}{ 2}}}\implies 3^{\color{red}{ 6}}x^{\frac{7}{2}\cdot {\color{red}{ 6}}+2\cdot {\color{red}{ 2}}}\)

OpenStudy (anonymous):

so... 729 times x^21+4? am I on the right track?

OpenStudy (jdoe0001):

so... 729 times x^21+4? am I on the right track? \(\large \checkmark\) notice that, when in a parentheses, you distribute the exponent and when you have the same base, but different exponent you keep the base, SUM the exponents

OpenStudy (anonymous):

Okay 729x^21+4? That cant be right... O.o

OpenStudy (jdoe0001):

well, heheh, it's if it's \(\Large \bf (3x^{\frac{7}{2}})^{\color{red}{ 6}}(x^2)^{\color{red}{ 2}}\)

OpenStudy (jdoe0001):

unless you mistyped it or missed something

OpenStudy (anonymous):

shouldn't the ^2 be times by 6 instead of 2? cause none of my answers say 729x^25

OpenStudy (jdoe0001):

ahemm no, because the ^2 is an exponent of ANOTHER component, while the ^6 is elsewhere they only merge because the BASE is the same

OpenStudy (anonymous):

but in the beginning of the problem it says (x^2)^6. im not trying to be a pain I just really need to figure this out... :/

OpenStudy (jdoe0001):

hmmm that's not what you have originally \(\huge \bf (3x^{\frac{7}{2}})^{\color{red}{ 6}}(x^2)^{\color{red}{ 2}}\)

OpenStudy (anonymous):

(3x^7/2)^6(x^2)^6 see that's our mix up

OpenStudy (jdoe0001):

\(\huge (3x^{^{\frac{7}{2}})^{\color{red}{ 6}}}(x^2)^{\color{red}{ 2}}\) tis what I read

OpenStudy (jdoe0001):

well. same procedure, do the same, distribute the exponent same bases, keep the base, add the exponents

OpenStudy (anonymous):

okay so the answer would be 729x^33?

OpenStudy (jdoe0001):

okay so the answer would be 729x^33? \(\large \checkmark\)

OpenStudy (anonymous):

Okay thank you!

OpenStudy (jdoe0001):

yw